The rate constant of a chemical reaction increased from 0.100 s−1 to 3.00 s−1 upon raising the temperature from 25.0 ∘C to 37.0 ∘C .May I please ask what is the activation energy of the reaction?Thanks
1
Expert's answer
2020-06-11T10:31:55-0400

T1=25+273=298 K

T2= 37+273=310 K

Ea=2.303*RT1T2*lg(k2/k1)/(T2-T1)=2.303*8.314*298*310*1.477/12=217711.4 J/mol=2.18*10^2 kJ/mol